OpenStudy (anonymous):

Solve: 6(x + 4) = 8(x − 2)

OpenStudy (uri):

xoya staaph

OpenStudy (uri):

I'll do this one

OpenStudy (anonymous):

Okay

OpenStudy (uri):

6x+24=8x-16 24+16=8x-6x 40=2x x=20

OpenStudy (uri):

40=2x* x=40/2 x=20